The square of the sum of a number and 3 is 16. Find the number. (There are two solutions.)
Answer by checkley77(12844)

(x+3)^2=16
x^2+6x+9-16=0
x^2+6x-7=0
(x+7)(x-1)=0
x+7=0
x=-7 answer.
x-1=0
x=1 answer.